# Automated Test Transfer Across Android Apps Using Large Language Models

### v1

@@ Abstract.
ccs: Software and its engineering Software notations and tools**footnotetext: These authors contributed equally to this work
@@ 1. Introduction
Testing mobile apps’ user interfaces (UIs) is crucial and ensures a seamless user experience across different devices and platforms. Manually testing mobile UIs requires significant time and effort and is prone to human error. Many approaches have been proposed in recent years to address these issues and automate the testing process (Li and Wu, 2006; Kaasila et al., 2012; Su et al., 2017a; Mao et al., 2016a; Amalfitano et al., 2012; Gu et al., 2019; Dong et al., 2020; Wang et al., 2020; Hu et al., 2018; Mon, [n. d.]; Fir, [n. d.]; Amalfitano et al., 2015; Anand et al., 2012; Choi et al., 2013; Hao et al., 2014; Koroglu et al., 2018; Jensen et al., 2013; Linares-Vásquez et al., 2015; Machiry et al., 2013; Mahmood et al., 2014; Mao et al., 2016b; Mirzaei et al., 2015; Yang et al., 2013; Zaeem et al., 2014; Zhang and Rountev, 2017; Su et al., 2017b; Liu et al., 2017; Ermuth and Pradel, 2016). Many of these automated testing approaches focus on crash detection or maximizing certain criterion such as activity coverage (Mao et al., 2016a; Wang et al., 2020; Su et al., 2017a; Moran et al., 2016; Gu et al., 2019; Dong et al., 2020). However, recent studies (Linares-Vásquez et al., 2017; Kochhar et al., 2015; Haas et al., 2021) have indicated that developers prefer UI tests that target specific functionalities of an app. For instance, in a shopping app, the tester wants to ensure that a user is able to successfully register, search for a product, and add the product to the shopping cart.
@@
Test Transfer—also known as Test Migration —is a solution proposed in recent years for automatically creating new usage-based tests for a mobile app by using already existing tests from a similar app (Lin et al., 2019a; Behrang and Orso, pear). The concept of test transfer is based on the idea that apps within a specific domain, such as shopping, mail clients, or web browsers, despite potential differences in appearance and the way they are programmed, share very similar functionalities. Therefore, it is possible to use the already existing usage-based tests from one app to automatically create tests that exercise analogous functionalities in another app within the same domain. For instance, use a test for validating the login functionality in Geek , a shopping app, to automatically create tests for validating the login functionality in other popular shopping apps such as Yelp or Zalando (zal, [n. d.]).
@@
The existing test transfer solutions can be divided into two categories: (1) similarity-based approaches, which focus on matching events between the source and target apps using a distance metric (Behrang and Orso, 2019, 2018; Lin et al., 2019b; Mishra et al., 2023; Khalili et al., 2024; Liu et al., 2022; Mariani et al., 2021; Liu et al., 2024d), and (2) classification approaches that define the test transfer problem as a classification task (Hu et al., 2018).
@@
The similarity-based techniques have demonstrated high effectiveness in cases where the source and target apps are very similar. However, these approaches face significant limitations in more complex cases, especially where a straightforward one-to-one correspondence between the event sequences representing specific functionalities in the source and target apps does not exist (Liu et al., 2024d; Zhao et al., 2020). This limitation is mainly due to the fact that these approaches focus on directly mapping events between the source and target apps. As a result, they encounter limitations in adapting to cases where such direct mappings do not exist, which is often the case in real-world apps.
@@
The only existing machine learning-based method, AppFlow (Hu et al., 2018), has shown to be highly dependent on the categories of apps that are used for its training, which hampers its generalizability. Furthermore, this technique requires considerable manual effort to adapt to a new app category, further limiting its usefulness in practice.
@@
Recent advances in AI have led to the development of Large Language Models (LLMs), which have shown significant promise in automating various software engineering tasks such as code generation (Liu et al., 2024c), code summarization (Ahmed and Devanbu, 2022), and software testing (Wang et al., 2024). These models, trained on a large amount of data, are capable of accurately understanding the semantics of user interfaces, including screens and widgets (Liu et al., 2024a). Intrigued by these results, we set out to investigate the degree to which the rich semantics embedded in LLMs can be applied to advance the state-of-the-art in test transfer.
@@
LLMigrate’s transfer process consists of two main steps: (1) Source Abstraction, and (2) Test Migration. In the source abstraction phase, the source test is converted to a version represented in natural language, which we call the Abstract Source Test. This is achieved by executing the source test on the source app to extract the required features and consulting with the LLM to understand the semantics of the source test. During the second step, test migration, the abstract source test in natural language, which was generated in the first phase, is adapted to the target app. This is achieved by dynamically exploring the target app, extracting features at each step, and selecting the optimal actions with the guidance of the LLM.
We evaluated LLMigrate by applying it to transfer UI tests for apps in five different app categories. In each category, tests were transferred across different apps, resulting in a total of 120 transfers. Both the tests and the apps are reused from the dataset introduced by CraftDroid (Lin et al., 2019b). The transfer outcomes are evaluated based on the precision, recall, reduction, and success rate metrics, and the results were compared to those obtained from the existing techniques evaluated on the same benchmark (Lin et al., 2019b; Zhang et al., 2024b; Liu et al., 2024d). LLMigrate achieved a 97.5% success rate, reducing the manual effort required to write tests from scratch by 91.1%. This represents an improvement of 9.1% in success rate and 38.2% in effort reduction compared to the best-performing technique previously evaluated on the CraftDroid benchmark. Notably, LLMigrate attains these gains with an average transfer time of 247 seconds, surpassing the efficiency of prior methods.
The work presented in this paper makes the following contributions:
  • The first approach that leverages multimodal LLMs for UI test transfer across mobile apps. Our technique is an end-to-end solution that does not require access to the source code of the apps, ensuring accurate execution without human intervention.
  • An extensive evaluation of our technique on a set of diverse real-world apps.
  • A public repository (rep, [n. d.]) containing the artifacts, including the framework capable of migrating tests between apps and evaluation data that foster future research in the area of test transfer.
The remainder of this paper is organized as follows. Section 2 provides the used terminology and an example motivating this work. Section 3 discussed LLMigrate’s approach. Section 4 details the evaluation of LLMigrate on real-world apps. Section 5 presents the threats to validity, and Section 6 discusses related research. Finally, Section 7 concludes the paper.
@@ 2. Background and Terminology
Figure 1. Registration test in DODuae (dod, [n. d.]).
Figure 2. Registration test in Zalando (zal, [n. d.]).
@@ 3.1.2. Abstraction
In this step, the augmented test, the output of the previous step, is translated to a representation in natural language, called the Abstract Source Test. To achieve this, we use the LLM Agent module, which is responsible for creating prompts and communicating with the off-the-shelf LLMs. To create the prompts for each of the tasks in our approach that require communication with LLM, we have manually created a prompt template that includes the task definition and the necessary structure to introduce each of the input features. This prompt template definition is a one-time manual effort needed for each task, such as test abstraction. Once the template is defined, for every query instance, the LLM Agent integrates the variable input features into the prompt template.
The utilized off-the-shelf LLM accepts two types of prompts: 1) system prompt, which is a general prompt that familiarizes the agent with the task and provides a general context, and 2) user prompt, which is task-specific. Prompt 3.1.2 represents the user prompt template for the abstraction task. Note that, in all of the prompts demonstrated throughout the paper, the text presented in color green, represents the variable input features that are substituted in the prompt template for each specific query. For each task, we also provide a system prompt (not shown here) that specifies the context for the LLM. Due to space constraints, the complete prompt details are available in our publicly accessible repository (rep, [n. d.]).
@@
This test contains three oracles and evaluates the registration functionality of the ”mobilapp.opencart.
doduae” app. It begins by verifying the presence of the dashboard element (oracle), then navigates to the ”Me” section (GUI event), and confirms the presence of the register button (oracle). The test proceeds by clicking the ”Register” button (GUI event), entering ”sample@gmail.com” as the email address and ”samplepassword” as the password (GUI events), and clicking the ”Register” button again (GUI event). The final step, which completes the test, is an oracle that checks for the presence of the text ”sample@gmail.com” to confirm successful registration.
@@ 3.2. Test Migration
The complete set can be found in our implementation (rep, [n. d.]), and if needed, extended.
@@
The rationale for using multimodal inputs for the screen understanding task is that relying solely on the textual info obtained from the layout hierarchy can sometimes be insufficient, which may lead to suboptimal event selection. For instance, consider the example of adding an entry in a Todo app, demonstrated in Figure 5. In this case, two highlighted widgets exist with textual attributes fab_add and fab_add_reminder,
respectively. By only relying on textual information, it can be ambiguous to determine which of the two widgets should be used to add a new task, since the term ”reminder” is often used interchangeably with ”task” or ”to-do” in the domain of to-do apps. However, by incorporating the visual information from the captured screenshot, the screen analyzer agent is able to use the images and icons and conclude that the functionality of the second widget is to allow the user to add a reminder to a to-do item.
@@
”wait_until_element_presence” or ”wait_until_element_invisible”. If you return an oracle event you just need to include just event_type and action fields not anything else. […]
@@
The LLM Agent consults the LLM using Prompt 5 to create the initial optimal event, which is expected to be in the JSON format, including the event triple (action, event_type, widget) defined in Section 2. In order to mitigate the challenge imposed by the non-determinism of LLMs, we took two specific actions for all the action selection prompts. Firstly, we set the temperature parameter of the LLM to zero to reduce its exploration abilities and make its responses more predictable. Second, to reduce the impact of LLM hallucination, we used a majority algorithm to determine the final generated event. This involves querying the LLM n times repeatedly for each task as a zero-shot prompt, leading to the receipt of n responses in the form of JSON objects. In a JSON object, a key is a unique identifier for a piece of data, and its corresponding value is the data itself. To generate the final event, the LLM Agent includes only the keys that appear more than m times among the n received responses and sets the value of these keys to the most frequently occurring value for each included key. Note that m and n are configurable parameters that are set to 2 and 3, respectively, in our current evaluation.
@@
Scenarios 2 and 3 occur when an invalid event is created by LLM, which are easily identifiable by the UI Automator component. This is because executing these events on the target app results in raising an exception. In these cases, again, another iteration of the migration loop begins. However, since the previous iteration was unsuccessful, additional measures need to be taken by Explorer, our primary coordinator, to address the failure. These measures are fourfold. First, the Explorer component removes the previously generated event that resulted in an exception from the list of executed events, as it should not be included in the final transferred test. Second, the Explorer records this event as a dead-end event to ensure that it will not be mistakenly considered a valid event in future iterations. It is important to note that dead-end events are saved based on the current transfer step, as an event might be a dead-end in one step but not in another one. During step 5 of the migration loop, the Explorer always checks to ensure that a suggested event by the LLM Agent has not been previously identified as a dead-end event. Third, the Explorer backtracks one step in the execution to undo the consequences of the last invalid generated event. The backtrack is achieved by restarting the app and re-executing all recorded events except the last invalid one. Finally, the Explorer needs to inform the LLM Agent that the generated event in the previous iteration was invalid. This results in the LLM Agent utilizing a specific prompt, asking the off-the-shelf LLM to repair its previous attempt.
@@ 4.1. Experimental Setup
LLMigrate is designed to be platform-agnostic and, therefore, capable of transferring tests across various devices and platforms. Our current implementation focuses on transferring tests across Android apps. For our evaluation, we utilized the publicly available dataset introduced by the authors of CraftDroid (Lin et al., 2019b). As shown in Table 1, this dataset contains tests from five different app categories: Browser, To Do List, Shopping, Mail Client, and Tip Calculator. The advantages of utilizing this dataset are two-fold. First, it enables the evaluation of our developed approach within the context of real-world apps. Second, since this dataset has also been utilized as the benchmark for evaluating the other existing test transfer techniques, such as CraftDroid (Lin et al., 2019b), TEMdroid (Zhang et al., 2024b), TRASM (Liu et al., 2022), and TREADROID (Liu et al., 2024d), it enables us to effectively compare our technique against the state-of-the-art approaches across various dimensions such as accuracy, usefulness, and performance.
@@
For evaluating the efficacy of LLMigrate, we utilize the precision and recall metrics introduced and utilized by the existing research targeting test transfer (Lin et al., 2019b; Zhang et al., 2024b; Liu et al., 2024d). Similar to the definition utilized by the existing research, true positives (TP) are events in the transferred test that exist in the ground truth. False positives (FP) are events in the transferred test that do not exist in the ground truth. Finally, false negatives (FN) are events that exist in the ground truth but are not present in the transferred test.
@@
Table 3 presents a comparative analysis of the precision and recall metrics obtained by LLMigrate, CraftDroid (Lin et al., 2019b) and TREADROID (Liu et al., 2024d). Consistent with prior research, we categorized all events into two main types: GUI events and oracle events, with system events classified under GUI events. To ensure a fair and accurate comparison, we aimed to capture results achieved by the other techniques that were derived solely from the 19 subject apps that are currently functional from the CraftDroid dataset, as shown in Table 1. This required the detailed analysis of each transferred test for other techniques, similar to the process we used to analyze tests transferred by LLMigrate. This was possible for CraftDroid and TREADROID, as the artifacts containing the transferred tests are publicly available for CraftDroid, and we obtained the corresponding artifacts for TREADROID upon communicating with the authors. This was not possible for TEMdroid as the final transferred tests are not publicly available, and we were unable to obtain the research artifacts that would allow us to process it for a fair comparison, even after contacting the authors. Furthermore, we were unable to successfully run TEMdroid as the implementation of certain components was not publicly available. Consequently, for TEMdroid, we used the average metrics reported in their paper. In this case, although the datasets are not identical, the comparison remains informative since our dataset shares 82.6% (19 out of 23) of the apps. We have not included the metrics obtained by TRASM (Liu et al., 2022) and ATM (Behrang and Orso, 2019), as TREADROID previously benchmarked its obtained results against these techniques and demonstrated superior performance across all metrics (Liu et al., 2024d).
@@
As demonstrated in Table 3, LLMigrate was able to achieve a total average precision of 98% for GUI and 94% for oracle events. Similarly, LLMigrate achieved an average recall of 99% for GUI events and 94% for oracle events. These results indicate that LLMigrate outperforms all the existing techniques in both average precision and recall metrics for both GUI and oracle events in total.
@@
The usefulness of a test is defined by how helpful it is in reducing the manual effort for a human tester. To measure usefulness, we utilized the reduction metric, defined by previous research in this area (Zhao et al., 2020). This metric compares the manual effort required to write the ground truth test from scratch to the effort required to manually transform the transferred test to be identical to the ground truth test. The manual effort is defined as the Levenshtein distance (Levenshtein, 1966) between the sequence of events of the transferred and ground truth tests. The reduction metric is calculated using the following equation:
Reduction=#Ground Truth EventsManual Effort# Ground Truth $\text{Reduction}=\frac{\#\text{Ground Truth Events}-\text{Manual Effort}}{\text{\# Ground Truth Events}}$
@@
In categories such as shopping, which involve more complex test cases to transfer, previous research  (Lin et al., 2019b; Liu et al., 2024d) has shown a negative reduction metric, indicating that it can be more efficient to write tests from scratch rather than rely on transfer methods followed by extensive manual edits to the generated tests. This insight underscores the significance of metrics like reduction, which directly reflect the decrease in manual effort, as opposed to focusing solely on precision and recall of the transferred events. Since tools in this domain are fundamentally intended to transfer complete tests to minimize manual work for test engineers, reduction serves as a more relevant measure of a tool’s practical effectiveness.
@@
Figure 7 presents the successful transfer rate achieved by various techniques across different app categories and in total. On average, LLMigrate was able to achieve a total of 97.5% successful transfer rate across all the 120 transfers. This means that, although the transferred test may have additional extra steps not present in the ground truth or reach the objective via a path that was not identical to our specific manually defined ground truth, 97.5% of the transferred tests successfully met their objectives, including executing the required functionality and asserting the appropriate conditions using the transferred oracles. This shows that LLMigrate was able to outperform the existing technique with the highest successful transfer rate, TREADROID, by almost 10%. Again, for TEMdroid, we were unable to obtain the exact value for this metric, but the total average success rate reported in a subsequent work by the same authors (Zhang et al., 2024a) indicates a far inferior successful transfer rate of 53%, which is 44% lower than LLMigrate.
@@
As presented in Section 4, LLMigrate was able to achieve promising results on the CraftDroid dataset. While this dataset is the main benchmark utilized by the existing transfer techniques, it includes older apps, some of which are no longer functional or supported. As the mobile app development field is evolving rapidly and new workflows are being introduced daily to enhance the user experience, we conducted a study to: 1) assess the effectiveness of LLMigrate on more recent, popular real-world apps, and 2) make the scope of the evaluation of LLMigrate broader, specifically addressing the limitations in the availability of functional apps within certain app categories in CraftDroid dataset.
In this study, we utilized an additional widely used new app in the Browser, To Do List, and Mail Client categories and included two new popular apps in the Shopping category, which had the fewest functional apps in the CraftDroid dataset. In each category, we transferred the existing test from the CraftDroid dataset to the new apps. Table 4 demonstrates the precision, recall, success rate, and reduction metrics achieved by LLMigrate on the new set of apps.
As we migrated tests from the CraftDroid dataset, which is a relatively old dataset, to more recent apps, our tool demonstrated notable effectiveness, achieving an average success rate of 98%, an average reduction score of 97% and an average transfer time of 238 seconds. These results align closely with those presented in the context of experiments involved CraftDroid dataset, further validating LLMigrate’s capability to handle the evolving workflows of modern apps. This indicates that transferring tests from older apps is beneficial, as it significantly reduces the need for manually writing tests from scratch for new apps and proves our technique’s usefulness at the time of publication.

### v2

@@ Abstract.
copyright: acmlicenseddoi: 10.1145/3728975journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: ISSTAarticle: ISSTA098publicationmonth: 7ccs: Software and its engineering Software notations and tools**footnotetext: These authors contributed equally to this work
@@ 1. Introduction
Testing mobile apps’ user interfaces (UIs) is crucial and ensures a seamless user experience across different devices and platforms. Manually testing mobile UIs requires significant time and effort and is prone to human error. Many approaches have been proposed in recent years to address these issues and automate the testing process (Li and Wu, 2006; Kaasila et al., 2012; Su et al., 2017b; Mao et al., 2016a; Amalfitano et al., 2012; Gu et al., 2019; Dong et al., 2020; Wang et al., 2020; Hu et al., 2018; Team, 2023; Lab, 2024; Amalfitano et al., 2015; Anand et al., 2012; Choi et al., 2013; Hao et al., 2014; Koroglu et al., 2018; Jensen et al., 2013; Linares-Vásquez et al., 2015; Machiry et al., 2013; Mahmood et al., 2014; Mao et al., 2016b; Mirzaei et al., 2015; Yang et al., 2013; Zaeem et al., 2014; Zhang and Rountev, 2017; Su et al., 2017a; Liu et al., 2017; Ermuth and Pradel, 2016). Many of these automated testing approaches focus on crash detection or maximizing certain criterion such as activity coverage (Mao et al., 2016a; Wang et al., 2020; Su et al., 2017b; Moran et al., 2016; Gu et al., 2019; Dong et al., 2020). However, recent studies (Linares-Vásquez et al., 2017; Kochhar et al., 2015; Haas et al., 2021) have indicated that developers prefer UI tests that target specific functionalities of an app. For instance, in a shopping app, the tester wants to ensure that a user is able to successfully register, search for a product, and add the product to the shopping cart.
@@
Test Transfer—also known as Test Migration —is a solution proposed in recent years for automatically creating new usage-based tests for a mobile app by using already existing tests from a similar app (Lin et al., 2019a; Behrang and Orso, pear). The concept of test transfer is based on the idea that apps within a specific domain, such as shopping, mail clients, or web browsers, despite potential differences in appearance and the way they are programmed, share very similar functionalities. Therefore, it is possible to use the already existing usage-based tests from one app to automatically create tests that exercise analogous functionalities in another app within the same domain. For instance, use a test for validating the login functionality in Geek , a shopping app, to automatically create tests for validating the login functionality in other popular shopping apps such as Yelp or Zalando (zal, 2024).
@@
The existing test transfer solutions can be divided into two categories: (1) similarity-based approaches, which focus on matching events between the source and target apps using a distance metric (Behrang and Orso, 2019, 2018; Lin et al., 2019b; Mishra et al., 2023; Khalili et al., 2024; Liu et al., 2022; Mariani et al., 2021; Liu et al., 2024b), and (2) classification approaches that define the test transfer problem as a classification task (Hu et al., 2018).
@@
The similarity-based techniques have demonstrated high effectiveness in cases where the source and target apps are very similar. However, these approaches face significant limitations in more complex cases, especially where a straightforward one-to-one correspondence between the event sequences representing specific functionalities in the source and target apps does not exist (Liu et al., 2024b; Zhao et al., 2020). This limitation is mainly due to the fact that these approaches focus on directly mapping events between the source and target apps. As a result, they encounter limitations in adapting to cases where such direct mappings do not exist, which is often the case in real-world apps.
@@
The only existing machine learning-based method, AppFlow (Hu et al., 2018), has shown to be dependent on the categories of apps that are used for its training, hampering its generalizability. This technique also requires considerable manual effort to adapt to a new app category, further limiting its usefulness.
@@
Recent advances in AI have led to the development of Large Language Models (LLMs), which have shown significant promise in automating various software engineering tasks such as code generation (Liu et al., 2024d), code summarization (Ahmed and Devanbu, 2022), and software testing (Wang et al., 2024a). These models, trained on a large amount of data, are capable of accurately understanding the semantics of user interfaces, including screens and widgets (Liu et al., 2024a). Intrigued by these results, we set out to investigate the degree to which the rich semantics embedded in LLMs can be applied to advance the state-of-the-art in test transfer.
@@
LLMigrate’s transfer process consists of two main steps: (1) Source Abstraction, and (2) Test Migration. In the source abstraction phase, the source test is converted to a version represented in natural language, which we call the Abstract Source Test. This is achieved by executing the source test on the source app to extract the required features and consulting with the LLM to understand the semantics of the source test. During the second step, test migration, the abstract source test in natural language, which was generated in the first phase, is adapted to the target app. This is achieved by dynamically exploring the target app, extracting features at each step, and selecting the optimal actions with the help of LLM.
We evaluated LLMigrate by applying it to transfer UI tests for apps in five different app categories. In each category, tests were transferred across different apps, resulting in a total of 120 transfers. Both the tests and the apps are reused from the dataset introduced by CraftDroid (Lin et al., 2019b). The transfer outcomes are evaluated based on the precision, recall, reduction, and success rate metrics, and the results were compared to those obtained from the existing techniques evaluated on the same benchmark (Lin et al., 2019b; Zhang et al., 2024a; Liu et al., 2024b). LLMigrate achieved a 97.5% success rate, reducing the manual effort required to write tests from scratch by 91.1%. This represents an improvement of 9.1% in success rate and 38.2% in effort reduction compared to the best-performing technique previously evaluated on the CraftDroid benchmark. Notably, LLMigrate attains these gains with an average transfer time of 247 seconds, surpassing the efficiency of prior methods.
@@ 2. Background and Terminology
Figure 1. Registration test in DODuae (dod, 2024).
Figure 2. Registration test in Zalando (zal, 2024).
@@ 3.1.1. Test Augmentation
We support a limited set of UI interactions and primarily focus on two conditions for oracles: (1) the presence of an element, and (2) the invisibility of an element. The set of UI events and oracles supported in our work is consistent with those in the prior test transfer literature (Lin et al., 2019b; Zhang et al., 2024a; Liu et al., 2022, 2024b). Our augmentation module is available publicly on the project repository (rep, 2025).
@@ 3.1.2. Abstraction
In this step, the augmented test, the output of the previous step, is translated to a representation in natural language, called the Abstract Source Test, which serves as a short, one-paragraph summary of the source test that describes each of its UI and oracle events.
To achieve this, we use the LLM Agent module, which is responsible for creating prompts and communicating with the off-the-shelf LLMs. To create the prompts for each of the tasks in our approach that require communication with LLM, we have manually created a prompt template that includes the task definition and the necessary structure to introduce each of the input features. This prompt template definition is a one-time manual effort needed for each task, such as test abstraction. Once the template is defined, for every query instance, the LLM Agent integrates the variable input features into the prompt template.
The utilized off-the-shelf LLM accepts two types of prompts: 1) system prompt, which is a general prompt that familiarizes the agent with the task and provides a general context, and 2) user prompt, which is task-specific. Prompt 3.1.2 represents the user prompt template for the abstraction task, defining the requirements for the abstract test. This includes specifications such as its length and the information it should include. Note that, in all of the prompts demonstrated throughout the paper, the text presented in color green, represents the variable input features that are substituted in the prompt template for each specific query. For each task, we also provide a system prompt (not shown here) that specifies the context for the LLM. Due to space constraints, the complete prompt details are available in our publicly accessible repository (rep, 2025).
@@
This test contains three oracles and evaluates the registration functionality of the ”mobilapp.opencart.doduae” app. It begins by verifying the presence of the dashboard element (oracle), then navigates to the ”Me” section (GUI event), and confirms the presence of the register button (oracle). The test proceeds by clicking the ”Register” button (GUI event), entering ”sample@gmail.com” as the email address and ”samplepassword” as the password (GUI events), and clicking the ”Register” button again (GUI event). The final step, which completes the test, is an oracle that checks for the presence of the text ”sample@gmail.com” to confirm successful registration.
@@ 3.2. Test Migration
The complete set can be found in our implementation (rep, 2025), and if needed, extended.
@@
The rationale for using multimodal inputs for the screen understanding task is that relying solely on the textual info obtained from the layout hierarchy can sometimes be insufficient, which may lead to suboptimal event selection. For instance, consider the example of adding an entry in a Todo app, demonstrated in Figure 5. In this case, two highlighted widgets exist with textual attributes fab_add and fab_add_reminder, respectively. By only relying on textual information, it can be ambiguous to determine which of the two widgets should be used to add a new task, since the term ”reminder” is often used interchangeably with ”task” or ”to-do” in the domain of to-do apps. However, by incorporating the visual information from the captured screenshot, the screen analyzer agent is able to use the images and icons and conclude that the functionality of the second widget is to allow the user to add a reminder to a to-do item.
@@
”wait_until_element_presence” or ”wait_until_element_invisible”. If you return an oracle event you just need to include event_type and action fields not anything else. […]
@@
The LLM Agent consults the LLM using Prompt 5 to create the initial optimal event, which is expected to be in the JSON format, including the event triple (action, event_type, widget) defined in Section 2. Examples for LLM responses that contain an event are shown in Figure 4. In order to mitigate the challenge imposed by the non-determinism of LLMs, we took two specific actions for all the action selection prompts. Firstly, we set the temperature parameter of the LLM to zero to reduce its exploration abilities and make its responses more predictable. Second, to reduce the impact of LLM hallucination, we used a majority algorithm to determine the final generated event. This involves querying the LLM n times repeatedly for each task as a zero-shot prompt, leading to the receipt of n responses in the form of JSON objects. In a JSON object, a key is a unique identifier for a piece of data, and its corresponding value is the data itself. To generate the final event, the LLM Agent includes only the keys that appear more than m times among the n received responses and sets the value of these keys to the most frequently occurring value for each included key. Note that m and n are configurable parameters that are set to 2 and 3, respectively, in our current evaluation.
@@
Scenarios 2 and 3 occur when an invalid event is created by LLM, which are easily identifiable by the UI Automator component. This is because executing these events on the target app results in raising an exception. In these cases, again, another iteration of the migration loop begins. However, since the previous iteration was unsuccessful, additional measures need to be taken by Explorer, our primary coordinator, to address the failure. These measures are fourfold. First, the Explorer component removes the previously generated event that resulted in an exception from the list of executed events, as it should not be included in the final transferred test. Second, the Explorer records this event as a dead-end event to ensure that it will not be mistakenly considered a valid event in future iterations. It is important to note that dead-end events are saved based on the current transfer step, as an event might be a dead-end in one step but not in another one. During step 5 of the migration loop, the Explorer always checks to ensure that a suggested event by the LLM Agent has not been previously identified as a dead-end event. Third, the Explorer backtracks one step in the execution to undo the consequences of the last invalid generated event. The backtrack is achieved by restarting the app, clearing the cache, and re-executing all recorded events except the last invalid one. Finally, the Explorer needs to inform the LLM Agent that the generated event in the previous iteration was invalid. This results in the LLM Agent utilizing a specific prompt, asking the off-the-shelf LLM to repair its previous attempt.
@@ 4. Evaluation
  5. RQ5.
    How well does LLMigrate perform in transferring tests across apps that are not previously seen by LLMs?
@@ 4.1. Experimental Setup
LLMigrate is designed to be platform-agnostic and, therefore, capable of transferring tests across various devices and platforms. Our current implementation focuses on transferring tests across Android apps. For our evaluation, we utilized the publicly available dataset introduced by the authors of CraftDroid (Lin et al., 2019b). As shown in Table 1, this dataset contains tests from five different app categories: Browser, To Do List, Shopping, Mail Client, and Tip Calculator. The advantages of utilizing this dataset are two-fold. First, it enables the evaluation of our developed approach within the context of real-world apps. Second, since this dataset has also been utilized as the benchmark for evaluating the other existing test transfer techniques, such as CraftDroid (Lin et al., 2019b), TEMdroid (Zhang et al., 2024a), TRASM (Liu et al., 2022), and TREADROID (Liu et al., 2024b), it enables us to effectively compare our technique against the state-of-the-art approaches across various dimensions such as accuracy, usefulness, and performance.
@@
As discussed in Section 3, our approach has three adjustable parameters: 1) maximum wrong tries at the Same Step, 2) total number of runs for majority voting (n), and 3) threshold for majority voting to include a field (m). We empirically observed the best-performing values for all these parameters and set them to 3, 3, and 2, respectively, in our evaluation.
@@
For evaluating the efficacy of LLMigrate, we utilize the precision and recall metrics introduced and utilized by the existing research targeting test transfer (Lin et al., 2019b; Zhang et al., 2024a; Liu et al., 2024b). Similar to the definition utilized by the existing research, true positives (TP) are events in the transferred test that exist in the ground truth. False positives (FP) are events in the transferred test that do not exist in the ground truth. Finally, false negatives (FN) are events that exist in the ground truth but are not present in the transferred test.
@@
Table 3 presents a comparative analysis of the precision and recall metrics obtained by LLMigrate, CraftDroid (Lin et al., 2019b) and TREADROID (Liu et al., 2024b). Consistent with prior research, we categorized all events into two main types: GUI events and oracle events, with system events classified under GUI events. To ensure a fair and accurate comparison, we aimed to capture results achieved by the other techniques that were derived solely from the 19 subject apps that are currently functional from the CraftDroid dataset, as shown in Table 1. This required the detailed analysis of each transferred test for other techniques, similar to the process we used to analyze tests transferred by LLMigrate. This was possible for CraftDroid and TREADROID, as the artifacts containing the transferred tests are publicly available for CraftDroid, and we obtained the corresponding artifacts for TREADROID upon communicating with the authors. This was not possible for TEMdroid as the final transferred tests are not publicly available, and we were unable to obtain the research artifacts that would allow us to process it for a fair comparison, even after contacting the authors. Furthermore, we were unable to successfully run TEMdroid as the implementation of certain components was not publicly available. Consequently, for TEMdroid, we used the average metrics reported in their paper. In this case, although the datasets are not identical, the comparison remains informative since our dataset shares 82.6% (19 out of 23) of the apps. We have not included the metrics obtained by TRASM (Liu et al., 2022) and ATM (Behrang and Orso, 2019), as TREADROID previously benchmarked its obtained results against these techniques and demonstrated superior performance across all metrics (Liu et al., 2024b).
@@
As demonstrated in Table 3, LLMigrate was able to achieve a total average precision of 98% for GUI and 94% for oracle events. Similarly, LLMigrate achieved an average recall of 99% for GUI events and 94% for oracle events. These results indicate that LLMigrate outperforms all the existing techniques in both average precision and recall metrics for both GUI and oracle events in total. Note that, due to the varying lengths of different scenarios, all reported average metrics are calculated based on the total number of events within each category or across all transfers rather than based on the achieved metric for each individual migration. Furthermore, a detailed analysis of the results for each individual migration is presented in our publicly available repository (rep, 2025).
@@
The usefulness of a test is defined by how helpful it is in reducing the manual effort for a human tester. To measure usefulness, we utilized the reduction metric, defined by previous research in this area (Zhao et al., 2020). This metric compares the manual effort required to write the ground truth test from scratch to the effort required to manually transform the transferred test to be identical to the ground truth test. The manual effort is defined as the Levenshtein distance (Levenshtein, 1966) between the sequence of events of the transferred and ground truth tests. The reduction metric is calculated using the following equation: (# Ground Truth Events - Manual Effort) / (# Ground Truth Events).
@@
In categories such as shopping, which involve more complex test cases to transfer, previous research  (Lin et al., 2019b; Liu et al., 2024b) has shown a negative reduction metric, indicating that it can be more efficient to write tests from scratch rather than rely on transfer methods followed by extensive manual edits to the generated tests. This insight underscores the significance of metrics like reduction, which directly reflect the decrease in manual effort, as opposed to focusing solely on precision and recall of the transferred events. Since tools in this domain are fundamentally intended to transfer complete tests to minimize manual work for test engineers, reduction serves as a more relevant measure of a tool’s practical effectiveness.
@@
Figure 7 presents the successful transfer rate achieved by various techniques across different app categories and in total. On average, LLMigrate was able to achieve a total of 97.5% successful transfer rate across all the 120 transfers. This means that, although the transferred test may have additional extra steps not present in the ground truth or reach the objective via a path that was not identical to our specific manually defined ground truth, 97.5% of the transferred tests successfully met their objectives, including executing the required functionality and asserting the appropriate conditions using the transferred oracles. This shows that LLMigrate was able to outperform the existing technique with the highest successful transfer rate, TREADROID, by almost 10%. Again, for TEMdroid, we were unable to obtain the exact value for this metric, but the total average success rate reported in a subsequent work by the same authors (Zhang et al., 2024b) indicates a far inferior successful transfer rate of 53%, which is 44% lower than LLMigrate.
@@
As presented in Section 4, LLMigrate achieved strong results on the CraftDroid dataset, which is the primary benchmark used by existing test transfer techniques. However, since many apps in this dataset are outdated or no longer supported, and given the rapid evolution of mobile app development and workflows, we conducted a follow-up study to evaluate LLMigrate on more recent, widely used real-world apps. This study aimed to broaden the evaluation scope and address the dataset’s limitations, particularly in categories with fewer functional apps. We selected additional up-to-date apps in the Browser, To Do List, Mail Client and Shopping categories. In each case, we transferred existing CraftDroid tests to the new apps and measured precision, recall, success rate, and reduction, as shown in Table 4. LLMigrate demonstrated strong performance on this new set, achieving an average success rate of 98%, an average reduction score of 97%, and an average transfer time of 238 seconds. These findings validate LLMigrate’s effectiveness in adapting to modern app workflows and highlight the value of transferring tests from older apps to reduce manual testing effort in current app development.